OpenStudy (anonymous):

There is a positive correlation between the number of times the Striped Ground Cricket chirps per second and the temperature in degrees Fahrenheit. If a scatter plot is made with the number of chirps on the horizontal axis and the trend line is found to be y = 3x + 25, then what would you predict the temperature in degrees Fahrenheit to be when there are 20 chirps per second? a. -1.6 b. 85 c. 135

OpenStudy (anonymous):

i think a? not sure.

jimthompson5910 (jim_thompson5910):

hint: x = 20 in this case

OpenStudy (anonymous):

B!

jimthompson5910 (jim_thompson5910):

oh wait, they want the temp

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

sry i jumped the gun a bit

OpenStudy (anonymous):

its alrite.

jimthompson5910 (jim_thompson5910):

y = 3x + 25 20 = 3x + 25 ... Plug in y = 20 instead 20-25 = 3x -5 = 3x -5/3 = x -1.6667 = x x = -1.6667 so the closest answer is -1.6

jimthompson5910 (jim_thompson5910):

so your first answer was correct

OpenStudy (anonymous):

heehee. thanks for checking!

jimthompson5910 (jim_thompson5910):

you're welcome, went a bit too fast, but good thing there's a way to verify this lol

OpenStudy (anonymous):

yep.:)

jimthompson5910 (jim_thompson5910):

and i can see that they put that b) 85 in there to throw people off...very tricky

jimthompson5910 (jim_thompson5910):

but I guess that's normal